# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v3

@@ Abstract
The deployment of extremely large antenna arrays (ELAAs) in wideband extremely large-scale multiple-input multiple-output (XL-MIMO) systems introduces significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties. When combined with the dual-wideband effects inherent to wideband systems, these phenomena fundamentally reshape the channel’s sparsity patterns in the angular-delay domain, rendering existing sparsity-based channel estimation methods inadequate. To address these challenges, this paper reconsiders the channel estimation problem for wideband XL-MIMO systems. By leveraging the spatial-chirp property of array responses, we quantitatively characterize the angular-delay domain sparsity of wideband XL-MIMO channels, revealing both global block sparsity and local common-delay sparsity. To effectively capture this structured sparsity, we propose a novel column-wise hierarchical prior model that incorporates precision parameter sharing mechanism and Markov random field (MRF) structure. Based on this prior model, the channel estimation task is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem. Tailored to the complex factor graph induced by this hierarchical prior, we develop an MMV-based hybrid message passing (MMV-HMP) algorithm. The algorithm performs message updates along the factor graph edges by selectively applying either the sum-product algorithm (SPA) or variational message passing (VMP) rules, depending on the factor-node structure and message tractability. Simulation results validate the effectiveness of the proposed column-wise hierarchical prior model through comprehensive ablation studies and demonstrate that the MMV-HMP algorithm consistently outperforms existing baselines, achieving robust estimation performance with moderate computational complexity.
@@
Wideband XL-MIMO systems, dual-wideband effects, near-field effects, channel estimation, Bayesian inference.
@@ I Introduction
Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2, 3]. By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[4, 5]. It also delivers substantial beamforming gains, effectively mitigating the severe path loss characteristic of millimeter-wave (mmWave) frequencies. Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [6, 7]. However, realizing these performance benefits critically depends on obtaining accurate channel state information (CSI), underscoring the importance of precise channel knowledge.
Typically, wideband XL-MIMO systems adopt orthogonal frequency division multiplexing (OFDM) to combat the frequency-selective fading of wireless channels [8]. However, due to the pronounced impact of dual-wideband effects, it is challenging to acquire accurate CSI. Specifically, the large system bandwidth shortens the OFDM symbol duration, potentially causing the propagation delay difference across the antenna array to exceed the symbol duration, leading to the spatial wideband effect [9]. Concurrently, the varying central frequencies of subcarriers introduce the frequency-wideband effect, where phase shifts differ across subcarrier channels, further complicating the channel estimation process.
Moreover, with the deployment of ELAAs and operation at higher frequency bands, near-field effects, such as the spherical wavefront effect [10, 11, 12, 13] and spatially non-stationary (SnS) properties [14, 15], become prominent, exacerbating the dual-wideband effects. For instance, the sparsity structure of SnS dual-wideband XL-MIMO channels in the angular-delay domain undergoes significant changes. Specifically, the curvature of the spherical wavefront renders a single spatial frequency insufficient to characterize a near-field propagation path, requiring multiple spatial frequencies instead. This spatial frequency spread, combined with the frequency-wideband effect, amplifies angular spread compared to conventional massive MIMO systems [9, 16]. Additionally, the spherical wavefront effect disrupts the linear path delay variations across adjacent antennas, while SnS properties influence both spatial frequency and delay spreads. These combined factors reshape the sparse patterns of SnS dual-wideband XL-MIMO channels in the angular-delay domain, significantly complicating channel estimation in wideband XL-MIMO systems.
@@ I-B Motivations and Contributions
Due to the randomness of the environment and user locations, spherical wavefront effects and SnS properties are inevitable in wideband XL-MIMO systems. However, the channel estimation problem that jointly considers the spherical wavefront effect, SnS properties, and dual-wideband effects has not been well addressed to date. Specifically, a well-established channel model for wideband XL-MIMO systems is lacking, and the sparsity patterns of these channels in the angular-delay domain have yet to be fully explored. Additionally, with the increasing number of antennas and subcarriers in XL-MIMO systems, there is a pressing need for computationally efficient channel estimation algorithms to handle the growing complexity of these systems. To address these issues, this paper investigates channel estimation for XL-MIMO systems by incorporating the spherical wavefront effect, SnS properties, and dual-wideband effects. The main contributions of this paper are summarized as follows:
@@
  • Exploiting the sparsity in the angular-delay domain, the SnS dual-wideband XL-MIMO channel estimation problem is formulated as a multiple measurement vector (MMV)-based sparse recovery task. Recognizing the importance of an accurate prior model in channel reconstruction, we introduce a novel column-wise hierarchical prior, which incorporates both a precision parameter sharing mechanism and a Markov random field (MRF) structure. This design enables the model to effectively capture global block sparsity and local common-delay sparsity, thereby providing a more structured and accurate representation of the underlying channel characteristics.
  • To achieve efficient Bayesian inference on the intricate factor graph, resulting from the incorporation of both the precision parameter sharing mechanism and the MRF structure, we propose a MMV-based hybrid message passing (MMV-HMP) algorithm that jointly leverages belief propagation (BP) and variational message passing (VMP) rules. Thanks to this flexible message update strategy, the MMV-HMP algorithm can effectively manage the message updates for latent precision parameters with non-Gaussian distributions and accommodate the loops inherent in the complex factor graph, thereby enabling accurate and reliable inference.
  • Compared to existing approaches, the proposed MMV-HMP algorithm achieves superior performance by jointly capturing the global block sparsity and local common-delay sparsity inherent in SnS dual-wideband XL-MIMO channels. This enables robust channel estimation while maintaining moderate computational complexity. In addition, the effectiveness of the proposed column-wise hierarchical prior model is validated through comprehensive ablation studies.
@@ I-C Organization and Notations
The rest of the paper is organized as follows: Section II introduces the mmWave XL-MIMO system model. Section III develops a spatial-frequency channel model, incorporating spherical wavefront effects, SnS properties, and dual-wideband effects, and analyzes the angular-delay characteristics of SnS dual-wideband XL-MIMO channels. Section IV formulates the channel estimation problem as a sparse signal recovery task and proposes a column-wise hierarchical sparse prior model. Section V presents the MMV-HMP algorithm. Simulation results and conclusions are provided in Sections VI and VII, respectively.
Notations: lower-case letters, bold-face lower-case letters, and bold-face upper-case letters are used for scalars, vectors and matrices, respectively; The superscripts $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ stand for transpose and conjugate transpose, respectively; $\odot$ denotes the Hadamard product; $\mathbf{I}_{N}$ denotes a $N\times N$ identity matrix; $\mathbb{C}^{M\times N}$ denotes a $M\times N$ complex matrix. $\mathbf{1}_{N}$ denotes a $N\times 1$ all-one vector. In addition, a random variable $x\in\mathbb{C}$ drawn from the complex Gaussian distribution with mean $x_{0}$ and variance $v$ is characterized by the probability density function $\mathcal{CN}(x;m,v)=\exp\{-{\left|x-m\right|^{2}}/{v}\}/{\pi v}$; a random variable $\gamma\in\mathbb{R}$ from Gamma distribution with mean $a/b$ and variance $a/b^{2}$ is characterized by the probability density function $\mathcal{G}a(\gamma;a,b)\propto\gamma^{a-1}\exp(-\gamma b)$.
@@ II System Model
Consider a mmWave XL-MIMO OFDM system, as shown in Fig. 1, where a base station (BS) utilizes an ELAA comprising $N$ antennas arranged in a uniform linear configuration. To reduce hardware cost and energy consumption, the BS adopts a hybrid analog-digital architecture, where $N_{\mathrm{RF}}$ radio frequency (RF) chains are connected to the ELAA through phase shifters. The spacing between adjacent antennas is $d=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$, with $c$ and $f_{c}$ indicating the speed of light and the central carrier frequency, respectively. For the uplink channel estimation, we assume the $J$ users transmit mutual orthogonal pilot sequences to the BS [30, 26, 31], e.g., orthogonal time or frequency resources are utilized for different users to transmit pilot sequences. Therefore, channel estimation for each user is independent. Without loss of generality, we consider an arbitrary user and assume $K$ subcarriers are uniformly selected from the available subcarriers to carry pilot symbols.
Figure 1: Illustration of the XL-MIMO system model.
@@ III-A Channel Modeling for SnS XL-MIMO Systems
Assume that there are $L$ propagation paths between the BS and user. Denote =$\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l\geq 2$ indicates reflection or diffraction path, as illustrated in Fig. 1. Denote $\tau_{l,n}$ as the time delay of the $l$-th path for the $n$-th antenna. Then the time-invariant baseband channel impulse response of the $n$-th received antenna can be given by[32]
@@
where $\boldsymbol{\phi}_{l}=[n_{l,s},n_{l,e}]$ denotes the VR of the $l$-th path with $n_{l,s}$ and $n_{l,e}$ indicating the start and end index [^1] For the diffraction paths, $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ is associated with the diffraction coefficient. As such, $\kappa_{l,n}$ can be set within $10<\kappa_{l,n}<$ 1, with the reference value chosen as the element with the highest received power. Thus, we assume $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ obeys a uniform distribution between 0 and 1. Additionally, we assume continuous VRs for simplicity. Notably, the proposed method is not restricted to continuous VRs and can be extended to scenarios involving disjoint VRs, where the visible antennas may be distributed across multiple disjoint regions for a specific path.. Assume the distance and direction of the $l$-th path to the reference antenna element as $r_{l}$ and $\vartheta_{l}$, as illustrated in Fig. 1. Then, $\tau_{l,n}$, for any $l\in\mathcal{L}$, can be rewritten as
@@
$\overset{(a)}{\approx}$ $r_{l}-nd\cos\vartheta_{l}+\frac{n^{2}d^{2}}{2r_{l}}{\sin^{2}\vartheta_{l}}.$
where $(a)$ is obtained by using the Fresnel approximation, i.e., $\sqrt{1+x}\approx 1+x/2-x^{2}/8$. In this Taylor expansion, the current point is set at $n=0$, and terms of third order and higher are neglected. Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}-n{\psi_{l}}/{f_{c}}+n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$. According to (6) and (7), the channel impulse response in (4) can be further expressed as
@@
where $\alpha_{l}\triangleq\tilde{\alpha}_{l}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l}}$ is the equivalent complex path gain. Applying the Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (9)[33, 34], as shown in the top of this page.
@@
Further, Eq. (9) can be expressed in a more compact form as
@@
Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each pilot subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K-1$. The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in Eq. (10) with $f$ replaced by $f_{k}$. Thus, the overall spatial-frequency channel matrix $\mathbf{H}$ can be reformulated as
@@
Eq. (13) presents a comprehensive channel model for XL-MIMO systems, integrating spherical wavefront effects, SnS properties, as well as spatial and frequency-wideband characteristics. Specifically, the spatial-domain steering vector $\mathbf{b}(\psi_{l},\varphi_{l})$ is coupled with the SnS indicator vector $\boldsymbol{\kappa}_{l}$, the frequency response vector $\mathbf{a}(\tau_{l})$, and the frequency-dependent phase matrix $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$. However, due to the high dimensionality of spatial-frequency channels in mmWave XL-MIMO systems, direct estimation becomes computationally prohibitive. Thus, it is essential to leverage channel sparsity in a suitable transform domain.

### v4

@@ Abstract
The deployment of extremely large antenna arrays (ELAAs) in extremely large-scale multiple-input multiple-output (XL-MIMO) systems introduces significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties. When combined with the dual-wideband effects inherent to wideband systems, these phenomena fundamentally alter the channel’s sparsity patterns in the angular-delay domain, rendering existing estimation methods insufficient. To address these challenges, this paper reconsiders the channel estimation problem for wideband XL-MIMO systems. Leveraging the spatial-chirp property of array responses, we first quantitatively characterize the angular-delay domain sparsity of wideband XL-MIMO channels, revealing both global block sparsity and local common-delay sparsity. To effectively capture this structured sparsity, we then propose a novel column-wise hierarchical prior model that integrates a precision sharing mechanism and a Markov random field (MRF) structure. Building on this prior model, the channel estimation task is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem. Tailored to the complex factor graph induced by this hierarchical prior, we develop a MMV-based hybrid message passing (MMV-HMP) algorithm. This algorithm performs message updates along the edges of the factor graph, and selectively applies either the variational message passing (VMP) or sum-product (SP) rules, depending on the factor-node structure and message tractability. Simulation results validate the effectiveness of the proposed column-wise hierarchical prior model through ablation studies and demonstrate that the MMV-HMP algorithm, while maintaining moderate computational complexity, consistently outperforms existing baselines which fail to capture the structured sparsity of wideband XL-MIMO channels.
@@
Wideband XL-MIMO systems, dual-wideband effects, near-field effects, channel estimation, message passing.
@@ I Introduction
Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2]. By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[3, 4, 5]. It also delivers substantial beamforming gains, effectively compensating the severe path loss of millimeter-wave (mmWave) frequencies. Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [6, 7]. However, realizing these performance benefits critically depends on the accurate channel state information (CSI).
Typically, wideband XL-MIMO systems adopt orthogonal frequency division multiplexing (OFDM) to combat the frequency-selective fading of wireless channels [8]. However, due to the pronounced impact of dual-wideband effects, it is challenging to acquire accurate CSI. Specifically, the large system bandwidth shortens the OFDM symbol duration, potentially causing the propagation delay difference across the antenna array to exceed the symbol duration, leading to the spatial wideband effect [9]. Concurrently, the varying central frequencies of subcarriers introduce the frequency-wideband effect, where phase shifts differ across subcarrier channels, further complicating the acquisition of CSI.
Moreover, with the deployment of ELAAs and operation at higher frequency bands, near-field effects, such as the spherical wavefront effect [10, 11, 12] and spatially non-stationary (SnS) properties [13, 14], become prominent, exacerbating the dual-wideband effects. For instance, the sparsity structure of XL-MIMO channels in the angular-delay domain alters significantly. Specifically, the curvature of the spherical wavefront means a single spatial frequency is no longer sufficient to characterize a propagation path; instead, multiple spatial frequencies are required. This spatial frequency spread, combined with the frequency-wideband effect, results in a more pronounced angular spread compared to that of conventional massive MIMO systems [9, 15]. Additionally, the spherical wavefront effect disrupts the linear variations of path delay across antennas, while SnS properties influence both spatial frequency and delay spreads. These combined factors reshape the sparse patterns of XL-MIMO channels in the angular-delay domain, significantly complicating channel estimation in wideband XL-MIMO systems.
@@ I-B Motivations and Contributions
Due to the randomness of the environment and user locations, spherical wavefront effects and SnS properties are inevitable in wideband XL-MIMO systems. However, the channel estimation problem that jointly considers the spherical wavefront effect, SnS properties, and dual-wideband effects has not been well addressed to date. Specifically, a well-established channel model for wideband XL-MIMO systems is lacking, and the sparsity patterns of these channels in the angular-delay domain have yet to be fully explored. Additionally, with the increasing number of antennas and subcarriers in XL-MIMO systems, there is a pressing need for computationally efficient channel estimation algorithms to handle the growing complexity. To address these issues, this paper investigates channel estimation for XL-MIMO systems by incorporating the spherical wavefront effect, SnS properties, and dual-wideband effects. The main contributions of this paper are summarized as follows:
@@
  • Exploiting the sparsity in the angular-delay domain, the channel estimation problem is formulated as a multiple measurement vector (MMV)-based Bayesian inference task. Recognizing the importance of an accurate prior model in channel reconstruction, we devise a novel column-wise hierarchical prior, which incorporates both a precision sharing mechanism and a Markov random field (MRF) structure. This design enables the model to effectively capture global block sparsity and local common-delay sparsity, thereby providing a more structured representation of the underlying channel characteristics.
  • To achieve efficient Bayesian inference, resulting from the hierarchical prior, we propose a MMV-based hybrid message passing (MMV-HMP) algorithm that jointly leverages sum-product (SP) and variational message passing (VMP) rules. Thanks to this flexible message update strategy, the MMV-HMP algorithm can effectively manage the message updates for latent precision parameters with non-Gaussian distributions and accommodate the loops inherent in the complex factor graph, thereby enabling accurate inference.
  • Compared to existing approaches, the proposed MMV-HMP algorithm achieves superior performance by jointly capturing the global block sparsity and local common-delay sparsity while maintaining moderate computational complexity. In addition, comprehensive ablation studies validate the effectiveness of the proposed column-wise hierarchical prior model, confirming its ability to capture the channel’s structured sparsity.
@@ I-C Organization and Notations
The rest of the paper is organized as follows: Section II introduces the mmWave XL-MIMO system model. Section III develops a spatial-frequency channel model, incorporating near-field and dual-wideband effects, and analyzes the angular-delay characteristics of XL-MIMO channels. Section IV formulates the channel estimation problem and proposes a column-wise hierarchical sparse prior model. Section V presents the MMV-HMP algorithm. Simulation results and conclusions are provided in Sections VI and VII, respectively.
Notations: lower-case letters, bold-face lower-case letters, and bold-face upper-case letters are used for scalars, vectors and matrices, respectively; The superscripts $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ stand for transpose and conjugate transpose, respectively; $\otimes$ denotes the Kronecker product; $\mathbf{I}_{N}$ denotes a $N\times N$ identity matrix; $\mathbb{C}^{M\times N}$ denotes a $M\times N$ complex matrix. $\mathbf{1}_{N}$ denotes a $N\times 1$ all-one vector. In addition, a random variable $x\in\mathbb{C}$ drawn from the complex Gaussian distribution with mean $x_{0}$ and variance $v$ is characterized by the probability density function $\mathcal{CN}(x;m,v)=\exp\{-{\left|x-m\right|^{2}}/{v}\}/{\pi v}$; a random variable $\gamma\in\mathbb{R}$ from Gamma distribution with mean $a/b$ and variance $a/b^{2}$ is characterized by the probability density function $\mathcal{G}a(\gamma;a,b)\propto\gamma^{a-1}\exp(-\gamma b)$.
Figure 1: Illustration of the XL-MIMO system model.
@@ II System Model
Consider a mmWave XL-MIMO OFDM system, as shown in Fig. 1, where a base station (BS) utilizes an ELAA comprising $N$ antennas arranged in a uniform linear configuration. To reduce hardware cost and energy consumption, the BS adopts a hybrid analog-digital architecture, where $N_{\mathrm{RF}}$ radio frequency (RF) chains are connected to the ELAA through phase shifters. The spacing between adjacent antennas is $d=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$, with $c$ and $f_{c}$ indicating the speed of light and the central carrier frequency, respectively. For the uplink channel estimation, we assume the $J$ users transmit mutual orthogonal pilot sequences to the BS [29, 25, 30], e.g., orthogonal time or frequency resources are utilized for different users to transmit pilot sequences. Therefore, channel estimation for each user is independent. Without loss of generality, we consider an arbitrary user and assume $K$ subcarriers are uniformly selected from the available subcarriers to carry pilots.
@@ III-A Channel Modeling for SnS XL-MIMO Systems
Assume that there are $L$ propagation paths between the BS and user. Denote $\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l\geq 2$ indicates reflection or diffraction path, as illustrated in Fig. 1. Denote $\tau_{l,n}$ as the time delay of the $l$-th path for the $n$-th antenna. Then the time-invariant baseband channel impulse response of the $n$-th received antenna can be given by[31]
@@
where $\boldsymbol{\phi}_{l}=[n_{l,s},n_{l,e}]$ denotes the VR of the $l$-th path with $n_{l,s}$ and $n_{l,e}$ indicating the start and end index [^1] For the diffraction paths, $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ is associated with the diffraction coefficient. As such, $\kappa_{l,n}$ can be set within $0<\kappa_{l,n}<1$, with the reference value chosen as the element with the highest received power. Thus, we assume $\kappa_{l,n}\in\boldsymbol{\phi}_{l}$ obeys a uniform distribution between 0 and 1. Additionally, we assume continuous VRs for simplicity. Notably, the proposed method is not restricted to continuous VRs and can be extended to scenarios involving disjoint VRs, where the visible antennas may be distributed across multiple disjoint regions for a specific path.. Assume the distance and direction of the $l$-th path to the reference antenna element as $r_{l}$ and $\vartheta_{l}$, as illustrated in Fig. 1. Then, $\tau_{l,n}$, for any $l\in\mathcal{L}$, can be rewritten as
@@
$\overset{(a)}{\approx}$ $r_{l}-nd\cos\vartheta_{l}+\frac{n^{2}d^{2}}{2r_{l}}{\sin^{2}\vartheta_{l}},$
where $(a)$ is obtained by applying the second-order Taylor approximation with the current point set at $n=0$ and terms of third order and higher neglected. Notably, this approximation maintains high accuracy when the distance between the BS and the user or scatterer exceeds the Fraunhofer distance [29]. Additionally, unlike the Euclidean-domain model in [32], the approximation enables the array response vector to align with a spatial chirp form, simplifying subsequent analysis.
Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}-n{\psi_{l}}/{f_{c}}+n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$. According to (6) and (7), the channel impulse response in (4) can be further expressed as
@@
where $\alpha_{l}\triangleq\tilde{\alpha}_{l}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l}}$ is the equivalent complex path gain. Applying the Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (9)[33, 34], as shown in the top of the next page.
@@
Further, (9) can be expressed in a more compact form as
@@
Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each pilot subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K-1$. The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in (10) with $f$ replaced by $f_{k}$. Thus, the overall spatial-frequency channel matrix $\mathbf{H}$ can be reformulated as
@@
Eq. (13) presents an approximated channel model for XL-MIMO systems, integrating spherical wavefront effects, SnS properties, as well as spatial and frequency-wideband characteristics. Specifically, the spatial-domain steering vector $\mathbf{b}(\psi_{l},\varphi_{l})$ is coupled with the SnS indicator vector $\boldsymbol{\kappa}_{l}$, the frequency response vector $\mathbf{a}(\tau_{l})$, and the frequency-dependent phase matrix $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$.